# Longteng Guo1,3, Jie Jiang1,3, Jing Liu1,3 1Zidongtaichu Foundation Model Research Center, CASIA 2State Key Laboratory of Multimodal Artificial Intelligence Systems, CASIA

### v1

@@ 2.2 Task-shared Electrode Graph
Electrode Representation Extraction.
Consider a multi-task dataset defined as $Y=\{y_{1},y_{2},\ldots,y_{M}\}$. Each sample $y_{j}$ belongs to $\mathbb{R}^{E_{j}\times T\times C}$, where $E_{j}$ denote the number of electrodes. As illustrated in Figure 3 (right), for each sample $y_{j}$, a learnable special token $c\in\mathbb{R}^{C}$ is broadcast across all $E_{j}$ electrodes and appended to the end of the temporal sequence:
$y_{j}^{\prime}=\left[y_{j}||\ c\cdot\mathbf{1}_{E_{j}\times 1}\right]\in\mathbb{R}^{E_{j}\times(T+1)\times C}$ (6)
Leveraging the unidirectional attention mechanism inherent to autoregressive models, these special tokens facilitate the integration of local information from individual electrodes to synthesize global representations. Specifically, $y_{i}^{\prime}$ are then processed by the pre-trained ETE. Notably, the parameters of ETE are frozen during this stage, functioning solely as a feature extraction backbone. Subsequently, electrode representations are derived from the positions of special tokens in the output of ETE:
$z_{j}=\text{ETE}(y_{j}^{\prime})[:,-1,:]\in\mathbb{R}^{E_{j}\times C}$ (7)
@@
Network Structure.
We initially construct a graph network in which each node represents an electrode utilized during the pre-training stage. The total number of nodes is denoted by $|R(X)|$. Benefiting from the comprehensive data coverage in the pre-training, these nodes include nearly all electrodes commonly employed, encompassing those found in $Y$. Each node is represented by a learnable vector of length $C$. They form a fully interconnected graph $\mathcal{G}\in\mathbb{R}^{|R(X)|\times C}$. For each $z_{j}$, the electrodes it comprises are mapped to a subgraph $\mathcal{G}_{j}$ within $\mathcal{G}$. Upon the introduction of $z_{j}$ into the network, only the nodes contained within the subgraph $\mathcal{G}_{j}$ are activated. This activation process involves updating the representations of these specific nodes by adding the corresponding elements from $z_{j}$:
@@
The function $\text{diag}(z_{j})$ converts $z_{j}$ into a diagonal matrix, facilitating targeted activations that only influence corresponding nodes within $\mathcal{G}_{j}$. The indicator matrix $\mathbf{I}_{\mathcal{G}_{j}}$ ensures that updates are confined to these nodes, leaving others unaffected. The updated graph $\mathcal{G}$ facilitates the flow and interaction of spatial information between electrodes through a graph attention mechanism (Veličković et al., 2017):
@@
$\alpha_{mn}=\frac{\beta_{mn}\cdot\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{T}[\mathbf{W}h_{m}\|\mathbf{W}h_{n}]\right)\right)}{\sum_{k\in\mathcal{N}(m)}\beta_{mk}\cdot\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{T}[\mathbf{W}h_{m}\|\mathbf{W}h_{k}]\right)\right)}$ (9)
@@
where $\alpha_{mn}$ represents the attention coefficient between nodes $m$ and $n$. $\mathbf{W}$ and $\mathbf{a}$ are the learnable mapping weights. $h_{m}$ is the representation of node $m$ in $\mathcal{G}$. We use $\mathcal{N}(m)$ to denote the neighbor set of node $m$. For the activated subgraph, a masking coefficient $\beta$ is introduced, where $\beta_{mn}$ equals 1 if both $m$ and $n$ are within it, and 0 otherwise. Based on the obtained attention coefficients, the interactions between nodes are as follows:
@@
$h_{m}^{\prime}=\sigma\left(\sum_{n\in\mathcal{N}(m)}\alpha_{mn}\mathbf{W}h_{n}\right)$ (10)
@@
where $\sigma$ represents the activation function (ReLU (Glorot et al., 2011) in this context). Similarly, the above operation is stacked across $K$ layers, with each layer employing a residual connection and pre-normalization. For various subgraphs $\mathcal{G}_{j}$ (i.e., different datasets or tasks) within the same batch, unified training is efficiently achieved by constructing corresponding mask matrices $\beta$. This approach allows the model to operate as a multi-task generalist. In terms of output, the graph network pools the representations of nodes within $\mathcal{G}_{j}$ and subsequently directs them to the relevant task-specific head for either classification or regression.
3 EXPERIMENTS
3.1 Experimental Settings
Model Variants.
Table 1: Configuration of EEGPT models.
Configuration Base Large Huge Giant
ETE Layers 3 9 12 20
TEG Layers 3 3 4 4
Head Size 32 32 64 64
Hidden Size 128 256 896 1,792
Attention Heads 4 8 14 28
Intermediate Size 512 1,024 3,584 7,168
Total Parameters 1.46M 11.29M 183.8M 1.09B
We have developed four architecture configurations of EEGPT: EEGPT-Base, EEGPT-Large, EEGPT-Huge, and EEGPT-Giant. The parameter counts for these models are as follows: EEGPT-Base is 1.46M, EEGPT-Large is 11.29M, EEGPT-Huge is 183.8M, and EEGPT-Giant is 1.09B. In the case of the ETE and TEG network, they share the same hidden size and number of attention heads. These increments, which approximately scale by an order of magnitude at each level, are achieved by expanding the depth and width of the network. For a more detailed analysis of scaling law, please refer to Sec 3.3.
Training Details.
We adopt AdamW (Loshchilov & Hutter, 2017) as the optimizer and conduct all training on 8 NVIDIA A800-SXM4-80G GPUs. To enhance training efficiency, we utilize DeepSpeed Zero Optimization Stage 2. During pre-training, all model scales are trained for 3 epochs using a consistent dataset of 37.5M samples, which collectively includes approximately 1B tokens. The batch size and learning rate are set to 4096 and 1e-4, respectively. The maximum duration for pre-training (for EEGPT-Giant) is capped at 20 hours. For multi-task fine-tuning, all model scales are trained for 10 epochs using a consistent dataset of 181K samples. The batch size and learning rate are maintained at 512 and 1e-4, respectively. The maximum training duration for multi-task learning (for EEGPT-Giant) is limited to 3 hours. In this stage, the pre-trained parameters of ETE are frozen, and only the newly introduced TEG network is actively trained. Please refer to the Appendix for further details.
Baseline Models.
The baseline models selected for comparison are divided into two distinct categories. The first category encompasses traditional and widely utilized architectures in the EEG domain, such as EEGNet (Lawhern et al., 2018), TSception (Ding et al., 2022), Conformer (Song et al., 2023), and LGGNet (Ding et al., 2023). These models are trained from scratch on the respective datasets without any pre-training. The second category includes cutting-edge pre-trained models, i.e., LaBraM (Jiang et al., 2024) and BIOT (Yang et al., 2024b), They are fine-tuned on the respective datasets using inherited pretrained parameters. More details regarding of the baseline models could be found in the Appendix.
Considering that no existing models have been evaluated on such a diverse range of downstream tasks, we have meticulously reproduced these models using their official code, hyperparameter configurations, and pretrained checkpoints. This reproduction aims to supplement the performance metrics for each task, facilitating a comprehensive comparison. It is important to note that all baseline results are derived from models fine-tuned for specific tasks, indicating that these are individual specialist models. In contrast, the results from EEGPT originate from a single generalist model.
Evaluation Details.
Table 2: Statistical analysis of 12 evaluation datasets. Our in-house data is denoted by †.
@@ 3.1 Experimental Settings
We evaluate our EEGPT using 12 datasets across 5 distinct tasks, as detailed in Table 2. For Emotion Recognition (ER), we utilize DEAP (Koelstra et al., 2011), FACED (Chen et al., 2023), SEED-IV (Zheng et al., 2018), and SEED-V (Liu et al., 2021). For Motor Imagery (MI) classification, we employ MIBCI (Cho et al., 2017) and BCI Competition IV-1 (Blankertz et al., 2007). For Mental Workload (MW) detection, we select EEGMat (Zyma et al., 2019) and STEW (Lim et al., 2018). For Sleeping Stage (SS) classification, we analyze data from EDF (Kemp et al., 2000) and HMC (Alvarez-Estevez & Rijsman, 2021). For Cross Modality (CM) tasks, we employed IMG, our proprietary dataset, and SPE (Nguyen et al., 2017). All datasets use accuracy as the performance metric. Further descriptions and processing details are available in the Appendix.
@@
Across all datasets, we adopt a cross-subject paradigm. Specifically, we partition each dataset in the multi-task set into training, validation, and test splits using an 8:1:1 ratio, ensuring no overlap of subjects among these splits. To minimize variability, we calculate the average accuracy and standard deviation from results obtained using five distinct random seeds.
@@
Table 3: Evaluation on EEG Benchmarks. The column "One Model?" indicates whether the results for these benchmarks originate from the same model. The results in bold and underline are the best and second-best results, respectively.
@@ 3.2 Performance Evaluation
Table 3 presents a performance comparison across 12 datasets, illustrating that EEGPT, despite being a generalist model, consistently surpasses specialist models that have been fine-tuned for specific tasks. Specifically, EEGPT-Giant achieves an average accuracy improvement of 5.07% on the ER task, 6.05% on the MI task, 8.50% on the MW task, 11.20% on the SS task, and 5.10% on the CM task compared to the best performances by these specialist models. Moreover, as the model scales, there is a clear and consistent upward trend in performance improvement.
@@
Interestingly, our findings reveal that specialist models with pre-training appear to perform slightly worse than those trained from scratch. One intuitive hypothesis is that current mainstream EEG pre-training models are often based on large-scale seizure data, which exhibits domain discrepancy from typical EEG data used in general downstream tasks. This mismatch likely hampers the efficacy of transfer learning. Nonetheless, EEGPT models demonstrate considerable versatility and effectiveness across a diverse array of tasks, thereby robustly validating its utility and performance.
3.3 Ablation Study
@@
In this section, we conduct a detailed ablation analysis of the proposed training recipe. It is important to note that the findings are consistent across models of four different scales. Due to space limitations, we uniformly present the numerical results based on EEGPT-Large.
Scaling law for model size preliminarily emerges.
@@
Figure 4: Scaling laws for model size: (a) Pre-training loss curves of EEGPT with varying parameter scales; (b) Performance of EEGPT on 5 downstream tasks across different parameter scales.
@@
Figure 4 (a) compares the convergence curves of the autoregressive reconstruction loss across Base, Large, Huge, and Giant models. The results indicate that as the number of model parameters increases, the fit to the pre-training data improves, which is directly reflected in the final converged loss values. It largely indicates that models of a larger scale have absorbed more prior knowledge. We analyze the average performance of the four models across five tasks, as shown in Figure 4 (b). For all tasks, a positive correlation between performance and model size is evident, suggesting that larger models can effectively transfer more pre-training knowledge to a wide range of downstream tasks. This represents the first effective exploration and validation of the scaling law for autoregressive models in the EEG domain. We believe that with further increases in training scale, autoregressive architectures may exhibit enhanced generalization and versatility for EEG analysis.
Scaling law for training data preliminarily emerges.
@@
In this section, we delve into another critical dimension: the scaling laws of training data. For our analysis, we randomly shuffle 1B tokens designated for pre-training and distribute them into five groups: 0B, 0.25B, 0.5B, 0.75B, and 1B tokens. Notably, the group with 0B tokens represents the absence of pre-training. We conduct pre-training across these varied data volumes. After freezing these pre-trained models, we perform multi-task fine-tuning, keeping training steps and settings consistent. The corresponding results are presented in Figure 5. As demonstrated in the figure, there are evident performance improvements across all five tasks as the volumes of pre-training data increase. These improvements are initially substantial but gradually taper off as data volumes expand. Similar patterns are also observed in the field of NLP (Kaplan et al., 2020). Given that the trend of performance improvement with increasing data has not yet diminished, we believe that by further expanding the dataset, EEGPT could achieve even better performance.
@@
Figure 5: Scaling laws for data volume: As the size of training data increases, performance improvements are observed consistently across 5 tasks.
Autoregression outperforms bidirectional masked pre-training.

### v2

@@ III-C Multi-task Transfer Learning
Consider a multi-task dataset $Y=\{y_{1},y_{2},\ldots,y_{M}\}$. Each sample $y_{j}$ belongs to $\mathbb{R}^{E_{j}\times T\times D}$, where $E_{j}$ denote the number of electrodes. As illustrated in Fig. 3 (right), for each sample $y_{j}$, a learnable special token $c\in\mathbb{R}^{D}$ is broadcast across all $E_{j}$ electrodes and appended to the end of the temporal sequence:
$y_{j}^{\prime}=\left[y_{j}||\ c\cdot\mathbf{1}_{E_{j}\times 1}\right]\in\mathbb{R}^{E_{j}\times(T+1)\times D}$ (8)
Leveraging the unidirectional attention mechanism inherent to autoregressive models, these special tokens facilitate the integration of local information to synthesize global representations for each electrode. $y_{i}^{\prime}$ is then processed by the pre-trained ETE. Notably, the parameters of ETE are frozen during this stage, functioning solely as a feature extraction backbone. Subsequently, electrode representations are derived from the positions of special tokens in the output of ETE:
$z_{j}=\left.\text{ETE}(y_{j}^{\prime})\right|_{t=T+1}\in\mathbb{R}^{E_{j}\times D}$ (9)
@@
We initially construct a graph network in which each node represents an electrode utilized during the pre-training stage. The total number of nodes is the total number of electrodes $E$ in Eq. (2). Benefiting from the comprehensive data coverage in the pre-training, these nodes include nearly all electrodes commonly employed, encompassing those found in $Y$. Each node is represented by a learnable vector of length $D$. They form a fully interconnected graph $\mathcal{G}\in\mathbb{R}^{E\times D}$. For each $z_{j}$, the electrodes it comprises are mapped to a subgraph $\mathcal{G}_{j}$ within $\mathcal{G}$. Upon the introduction of $z_{j}$ into the network, only the nodes contained within the subgraph $\mathcal{G}_{j}$ are activated. This activation process involves updating the representations of these specific nodes by adding the corresponding elements from $z_{j}$:
@@
The function $\text{diag}(z_{j})$ converts $z_{j}$ into a diagonal matrix, facilitating targeted activations that only influence corresponding nodes within $\mathcal{G}_{j}$. The indicator matrix $\mathbf{I}_{\mathcal{G}_{j}}$ ensures that updates are confined to these nodes, leaving others unaffected. The updated graph $\mathcal{G}$ facilitates the flow and interaction of spatial information between electrodes through a graph attention mechanism [28]:
$\alpha_{mn}=\text{ReLU}(\mathbf{a}^{T}[\mathbf{W}h_{m}\|\mathbf{W}h_{n}])$ (11)
Specifically, for a given node $m$, represented by $h_{m}\in\mathbb{R}^{1\times D}$, the learnable projection weights $\mathbf{W}$ and $\mathbf{a}$ are used to compute the relevance $\alpha_{mn}$ to another node $n$. The relevance scores are then normalized within all neighboring nodes $\mathcal{N}(m)$ of $m$ to obtain the attention coefficients:
@@
$\gamma_{mn}=\frac{\beta_{mn}\cdot\exp\left(\alpha_{mn}\right)}{\sum_{k\in\mathcal{N}(m)}\beta_{mk}\cdot\exp(\alpha_{mk})}$ (12)
@@
To ensure that interactions occur only within the activated subgraph $\mathbf{I}_{\mathcal{G}_{j}}$, a masking coefficient $\beta$ is introduced, where $\beta_{mn}=1$ if both nodes $m$ and $n$ are within the subgraph, and $\beta_{mn}=0$ otherwise. Based on the attention coefficients, the interactions between nodes are computed as follows:
@@
$h_{m}^{\prime}=\text{ReLU}\left(\sum_{n\in\mathcal{N}(m)}\alpha_{mn}\mathbf{W}h_{n}\right)$ (13)
@@
After interacting with nodes within the subgraph, the representation of node $m$ is updated to $h_{m}^{\prime}$. Similarly, the above operation is stacked across $K$ layers, with each layer employing a residual connection and pre-normalization. For various subgraphs $\mathcal{G}_{j}$ (i.e., different datasets or tasks) within the same batch, unified training is efficiently achieved by constructing corresponding mask matrices $\beta$. This approach allows the model to operate as a multi-task generalist. In terms of output, the graph network pools the representations of nodes within $\mathcal{G}_{j}$ and subsequently directs them to the relevant task-specific head for either classification or regression.
IV EXPERIMENTS
IV-A Datasets
We evaluate our BrainGPT using 12 datasets across 5 distinct tasks, as detailed in Table I. For Emotion Recognition (ER), we utilize DEAP [29], FACED [30], SEED-IV [31], and SEED-V [32]. For Motor Imagery (MI) classification, we employ MIBCI [33] and BCI Competition IV-1 [34]. For Mental Workload (MW) detection, we select EEGMat [35] and STEW [36]. For Sleeping Stage (SS) classification, we analyze data from EDF [37] and HMC [38]. For Cross Modality (CM) tasks, we employed IMG, and SPE [39].
Table I: Statistical analysis of 12 evaluation datasets.
@@ IV-A Datasets
For data preprocessing, all the EEG signals are resampled to 256 Hz. The signals are then filtered between 0.1 and 100 Hz and segmented into samples of four seconds. Each sample is further segmented into 25 tokens with an overlap rate of 0.875, while each token has 256 sampling points. Besides, we apply the z-score normalization. No further preprocessing or artifact correction methods are applied. Across all datasets, we adopt a cross-subject paradigm. Specifically, we partition each dataset in the multi-task set into training, validation, and test splits using an 8:1:1 ratio, ensuring no overlap of subjects among these splits. To minimize variability, we calculate the average accuracy and standard deviation from results obtained using five distinct random seeds.
IV-B Baseline Models
The baseline models selected for comparison are divided into two distinct categories. The first category encompasses traditional and widely utilized models in the EEG domain, such as EEGNet [40], TSception [41], Conformer [20], and LGGNet [42]. These models are trained from scratch on the respective datasets without pre-training. The second category includes cutting-edge pre-trained models, i.e., LaBraM [13], BIOT [14]. They are fine-tuned on the respective datasets using their publicly available pretrained weights. The detailed descriptions of the baseline models that we reproduce for comparison in this work are as follows:
  • EEGNet [40]: EEGNet is a compact convolutional neural network. It leverages depthwise and separable convolutions to facilitate feature extraction and classification.
  • TSception [41]: TSception is a multi-scale convolutional neural network, capable of learning discriminative representations across both time and channel dimensions. It incorporates a dynamic temporal layer to capture dynamic temporal and frequency representations, while an asymmetric spatial layer is employed to learn discriminative global and hemisphere representations.
  • Conformer [20]: Conformer is a compact convolutional Transformer model, capable of encapsulating both local and global features. It incorporates a convolutional module to learn low-level local features while employing a self-attention mechanism to extract global correlations within the local temporal features.
  • LGGNet [42]: LGGNet is a neurologically inspired graph neural network. It models the intricate relationships both within and between the brain’s functional regions.
  • BIOT [14]: BIOT is a self-supervised biosignal learning model that tokenizes biosignals of various formats into sentences. It can enable representation learning through a MAE-based pre-training strategy.
  • LaBraM [13]: LaBraM is a self-supervised EEG model, which adopts the vector-quantized neural spectrum prediction to train a neural tokenizer that encodes EEG patches into compact neural codes. Through MAE strategy, it can learn generic representations.
Considering that no existing models have been evaluated on such a diverse range of downstream tasks, we have meticulously reproduced these models using their official code, hyperparameter configurations, and pretrained checkpoints. This reproduction aims to supplement the performance metrics for each task, facilitating a comprehensive comparison. It is important to note that all baseline results are derived from models fine-tuned for specific tasks, indicating that these are individual specialist models. In contrast, the results from BrainGPT originate from a single generalist model.
IV-C Experimental Settings
IV-C1 Model Variants
We have developed four architecture configurations of BrainGPT: BrainGPT-Base, BrainGPT-Large, BrainGPT-Huge, and BrainGPT-Giant. The parameter counts for these models are as follows: BrainGPT-Base is 1.46M, BrainGPT-Large is 11.29M, BrainGPT-Huge is 183.8M, and BrainGPT-Giant is 1.09B. In the case of the ETE and TEG network, they share the same hidden size and number of attention heads. These increments, which approximately scale by an order of magnitude at each level, are achieved by expanding the depth and width of the network.
Table II: Configuration of BrainGPT models.
Configuration Base Large Huge Giant
ETE Layers 3 9 12 20
TEG Layers 3 3 4 4
Head Size 32 32 64 64
Hidden Size 128 256 896 1,792
Attention Heads 4 8 14 28
Intermediate Size 512 1,024 3,584 7,168
Total Parameters 1.46M 11.29M 183.8M 1.09B
IV-C2 Training Details
We adopt AdamW [43] as the optimizer and conduct all training on 8 NVIDIA A800-SXM4-80G GPUs. To enhance training efficiency, we utilize DeepSpeed Zero Optimization Stage 2. During pre-training, all model scales are trained for 3 epochs using a consistent dataset of 37.5M samples, which collectively includes approximately 1B tokens. The batch size and learning rate are set to 4096 and 1e-4, respectively. For multi-task fine-tuning, all model scales are trained for 10 epochs using a consistent dataset of 181K samples. The batch size and learning rate are maintained at 512 and 1e-4, respectively. In this stage, the pre-trained parameters of ETE are frozen, and only the newly introduced TEG network is actively trained. The detailed training protocols of BrainGPT are reported in Table III, which contains the specific hyper-parameter configurations for the Stage I: Autoregressive pre-training and the Stage II: Multi-task fine-tuning.
Table III: Training hyperparameters for BrainGPT of two stages.
Stage Hyperparameter Base Large Huge Giant
Stage I Lr 1e-4
Epoch 3.0
Precision BF16
Deepspeed Zero2 Zero2 Zero2 Zero3
LR Schedule cosine decay
Warmup Ratio 0.03
Batch Size per GPU 512
Gradient Checkpoint True
Stage II Lr 1e-4
Epoch 10
Precision BF16
Deepspeed Zero2 Zero2 Zero2 Zero3
LR Schedule cosine decay
Warmup Ratio 0.1
Batch Size per GPU 64
Gradient Checkpoint True
IV-D Performance Evaluation
@@
Table IV: Evaluation on EEG Benchmarks. The column ”One Model?” indicates whether the results for these benchmarks originate from the same model. The results in bold and underline are the best and second-best results, respectively.
@@
Table IV presents a performance comparison across 12 datasets, illustrating that BrainGPT, despite being a generalist model, consistently surpasses specialist models that have been fine-tuned for specific tasks. Specifically, BrainGPT-Giant achieves an average accuracy improvement of 5.07% on the ER task, 6.05% on the MI task, 8.50% on the MW task, 11.20% on the SS task, and 5.10% on the CM task compared to the best performances by these specialist models. Moreover, as the model scales, there is a clear and consistent upward trend in performance improvement.
@@
Interestingly, our findings reveal that specialist models with pre-training appear to perform slightly worse than those trained from scratch. One intuitive hypothesis is that current mainstream EEG pre-training models are often based on large-scale seizure data, which exhibits domain discrepancy from typical EEG data used in general downstream tasks. This mismatch likely hampers the efficacy of transfer learning. Nonetheless, BrainGPT models demonstrate considerable versatility and effectiveness across a diverse array of tasks, thereby robustly validating its utility and performance.
IV-E Ablation Study
@@
In this section, we conduct a detailed ablation analysis of the proposed training recipe. It is important to note that the findings are consistent across models of four different scales. Due to space limitations, we uniformly present the numerical results based on BrainGPT-Large.
IV-E1 Scaling law for model size preliminarily emerges
@@
Figure 4: Scaling laws for model size: (a) Pre-training loss curves of BrainGPT with varying parameter scales; (b) Performance of BrainGPT on 5 downstream tasks across different parameter scales.
@@
Fig. 4 (a) compares the convergence curves of the autoregressive reconstruction loss across Base, Large, Huge, and Giant models. The results indicate that as the number of model parameters increases, the fit to the pre-training data improves, which is directly reflected in the final converged loss values. It largely indicates that models of a larger scale have absorbed more prior knowledge. We analyze the average performance of the four models across five tasks, as shown in Fig. 4 (b). For all tasks, a positive correlation between performance and model size is evident, suggesting that larger models can effectively transfer more pre-training knowledge to a wide range of downstream tasks. This represents the first effective exploration and validation of the scaling law for autoregressive models in the EEG domain. We believe that with further increases in training scale, autoregressive architectures may exhibit enhanced generalization and versatility.
IV-E2 Scaling law for training data preliminarily emerges
@@
In this section, we delve into another critical dimension: the scaling laws of training data. For our analysis, we randomly shuffle 1B tokens designated for pre-training and distribute them into five groups: 0B, 0.25B, 0.5B, 0.75B, and 1B tokens. Notably, the group with 0B tokens represents the absence of pre-training. We conduct pre-training across these varied data volumes. After freezing these pre-trained models, we perform multi-task fine-tuning, keeping training steps and settings consistent. The corresponding results are presented in Fig. 5. As demonstrated in the figure, there are evident performance improvements across all five tasks as the volumes of pre-training data increase. These improvements are initially substantial but gradually taper off as data volumes expand. Similar patterns are also observed in the field of NLP [44]. Given that the trend of performance improvement with increasing data has not yet diminished, we believe that by further expanding the dataset, BrainGPT could achieve even better performance.
@@
Figure 5: Scaling laws for data volume: As the training data increases, performance improvements are observed consistently across 5 tasks.
IV-E3 Autoregression outperforms bidirectional masked pre-training
Current EEG pre-training models employ a masked signal reconstruction framework using bidirectional attention [13, 14]. In this framework, random segments of the signal are masked, and the resulting input is processed by an encoder that reconstructs these segments based on contextual information (i.e., MAE). To enable a rigorous comparison between MAE and autoregressive (AR) modeling, we conduct an in-depth analysis using three distinct reconstruction loss functions: $\ell_{1}$, $\ell_{2}$, and cosine. For fairness, both pre-training paradigms utilize the same model architecture and parameter settings. We report related results in Table V. For clarity, the standard deviations of the results presented have been omitted. The conclusions are twofold. First, for the three types of reconstruction loss, $\ell_{2}$ outperforms $\ell_{1}$, while $\cos$ shows the least efficacy. Second, irrespective of the distance metric employed, the AR architecture consistently outshines the MAE architecture, demonstrating a more than 2% average accuracy advantage. These findings align with and support existing research in the NLP domain [45, 46, 47]. Specifically, the unidirectional modeling task poses significant challenges, enabling the model to learn more robust representations. Additionally, AR effectively adapts to the temporal characteristics of EEG signals, capturing their patterns more directly and naturally.